Find the number of terms in the sequence 5,2,-1,...,-37
The next term in 3,6,12,24,48,...
The 100th term in 6,8,17,26,35,...
Given that the sequence 1,6,x,y,20,...is a Fibonacci sequence, find x and y.

Answer by stanbon(75887)

Find the number of terms in the sequence 5,2,-1,...,-37
The next term in 3,6,12,24,48,...
---
Note: It looks to me like the next term is -4
================================================
The 100th term in 6,8,17,26,35,...
a(100) = a(1)+(n-1)d
= 3 + 99*(-3)
= -294
===========================
Given that the sequence 1,6,x,y,20,...is a Fibonacci sequence, find x and y.
x = 1+6 = 7
y = 6+7 = 13
===================
Cheers,
Stan H.